Item 12 A current loop in a motor has an area of 1.75 cm .lt carries a 240 mA current in a uniform field of 0.62 T Part A What is the magnitude of the maximum torque on the current loop? Express your answer with the appropriate units. Value Units Submit My Answers Give Up Continue
Answer #1

magnetic moment u = I*A


magnetic field = B

torque = u x B = u*B**sintheta


for maximum torque theta = 90

maximum torque = u*B = I*A*B


maximum torque = 240*10^-3*1.75*10^-4*0.62 = 2.604*10^-5 Nm <<<<---answer